# Quench Spectroscopy for Dissipative and Non-Hermitian Quantum Lattice Models

### v1

@@ Abstract
We study the dynamics of the open Bose-Hubbard chain confined in the superfluid phase submitted to a sudden global quench on the dissipations and the repulsive interactions. The latter is investigated by calculating the equations of motion of relevant quadratic correlators permitting to study the equal-time connected one-body and density-density correlations functions. We then compute the quench spectral function associated to each observable to perform the quench spectroscopy of this dissipative quantum lattice model. This permits to unveil the quasiparticle dispersion relation of the Bose-Hubbard chain in the superfluid phase in the presence of loss processes. The applicability of the quench spectroscopy is also generalized to non-Hermitian quantum lattice models by considering the non-Hermitian transverse-field Ising chain in the paramagnetic phase.
1 Introduction
In the last decades, the major progress in the experimental control of quantum lattice models has given consequent momentum to the analytical and experimental investigation of the quench dynamics of isolated quantum systems [1, 2, 3, 4, 5, 6]. Recently, a new spectroscopy method has been introduced theoretically and verify numerically using tensor networks for isolated quantum spin and bosonic lattice models. The latter is based on the dynamical properties of the model induced by weak sudden global or local quenches [7, 8] and is referred to as quench spectroscopy. This new method to determine low-lying excitation spectra has been recently applied experimentally to the dipolar XY model using Rydberg atoms [9] and has provided promising results. However, since quantum systems cannot be purely isolated from their environment, it is natural that the study of open quantum systems has attracted a lot of interest in the last few years [10, 11, 12] by taking into account dissipative effects such as local dephasing noise [13, 14, 15, 16, 17], incoherent hopping [18] and local gain/loss processes  [19, 20, 21, 22, 23, 24, 25, 26, 27]. These open quantum lattice models are governed by the well-known Lindblad master equation  [28, 29, 30, 31] characterising the time evolution of the quantum system density matrix. However, the possibility to apply the quench spectroscopy for open quantum lattice models has not been investigated and corresponds to the main study of the present research work.
Experimentally, quantum simulators based on trapped ultracold atoms loaded in an artificial optical lattice generated by the interference of counter-propagating laser beams permit to simulate quantum lattice models [32, 33]. Using bosonic atoms, the Bose-Hubbard model can thus be engineered [34, 35]. Its out-of-equilibrium properties are accessible via quantum quenches realised experimentally by suddenly modifying the intensity of the laser beams controlling the artificial lattice depth and thus the ratio between the hopping amplitude and the two-body repulsive interaction strength [3, 36]. This tuning of the interaction ratio offers the possibility to scan the quantum phase transition between the gapless superfluid and Mott-insulating phases for commensurate fillings of the lattice. Loss processes, inducing a rich and new physics such as the quantum Zeno effect [37], the loss-assisted quantum control [38] or the loss-induced cooling effect [39], can be generated using this experimental platform based on ultracold bosonic atoms where different number of particles can be involved. For instance, two-body losses can be engineered by light-assisted inelastic two-body collisions which can also occur naturally in ultracold-atom experiments [40, 41, 42, 21, 43, 44].
In this work, we propose to investigate the dynamics of the open Bose-Hubbard chain driven out of equilibrium via sudden global quenches on the interactions as well as on the dissipations. The dissipative effects consist here in on-site two-body losses. The term on-site refers to a loss process involving a single lattice site. This specific quantum lattice model as well as its quench dynamics can be simulated experimentally using ultracold bosonic atoms [45, 3, 46, 47] including a controllable strength of the on-site two-body losses engineered by light-assisted inelastic collisions [42, 21, 43, 44]. The quench dynamics of the isolated Bose-Hubbard chain has already been extensively studied theoretically and numerically using for instance tensor network or variational Monte-Carlo techniques  [48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 7, 8]. Regarding the latter in the presence of dissipations, its quench dynamics induced by two-body losses has been under-researched analytically and numerically [59, 60, 61]. Most importantly, the possibility to perform quench spectroscopy to dissipative quantum lattice models has not been studied. The latter statement is also valid for non-Hermitian quantum lattice models. These exotic class of quantum systems appear naturally when studying the dynamics of dissipative quantum lattice models in the so-called no-click limit [31]. Consequently, we also propose here to investigate the quench spectroscopy of the non-Hermitian ($s=1/2$ short-range interacting) transverse-field Ising chain confined in its paramagnetic phase. This underlines the importance of the research work presented here together with the possibility to confirm our conclusions experimentally at least for generic dissipative quantum lattice models.
@@ 1 Introduction
We first introduce the theoretical approach based on the calculation of the equations of motion associated to the relevant quadratic bosonic correlators. The latter is valid in the superfluid-mean-field regime of the Bose-Hubbard chain for on-site two-body losses. The latter is benchmarked with a quasiparticle theory for the isolated Bose-Hubbard chain for a sudden global quench on the repulsive interactions. The former theory is adapted to investigate the double quench dynamics of the Bose-Hubbard chain where both the interactions and the dissipations are quenched suddenly and globally. This will permit to investigate the quench spectroscopy of the dissipative Bose-Hubbard chain. We then discuss the applicability of the quench spectroscopy method to non-Hermitian quantum lattice models by considering the quench dynamics of the non-Hermitian transverse-field Ising chain confined in the (paramagnetic) $z$ polarized phase.
@@
This article is organised as follows: in Sec 2, we start by introducing the model and the quench procedure. In Sec. 3, we benchmark the equations of motion approach with a quasiparticle approach. In Sec. 4, we discuss the dissipative and interaction quench dynamics of the Bose-Hubbard chain for on-site two-body losses. In Sec. 5, we investigate theoretically the quench spectroscopy of the latter dissipative quantum lattice model. In Sec. 6, we discuss the experimental feasibility as well as the limitations associated to the quench spectroscopy method. Then, in Sec. 7, the quench spectroscopy approach is extended to non-Hermitian quantum lattice models by applying the latter to the non-Hermitian transverse-field Ising chain confined in its paramagnetic phase. Finally in Sec. 8, we present our conclusions and possible extensions of the presented research work.
2 Model and quench procedure
@@
where $\hat{b}_{R}$ and $\hat{b}_{R}^{{\dagger}}$ correspond to the bosonic annihilation and creation operators acting on the lattice site $R\in\mathbb{N}$, $\hat{n}_{R}=\hat{b}^{{\dagger}}_{R}\hat{b}_{R}$ denotes the local occupation number operator associated to the lattice site index $R$, $J>0$ corresponds to the hopping amplitude, $U>0$ is the on-site repulsive two-body interaction strength. At equilibrium and zero-temperature, the quantum phase diagram of the BH chain has been extensively studied [62, 63]. The latter displays a gapless superfluid (SF) phase and a gapped Mott-insulating (MI) phase with the so-called Mott lobes, determined by the competition between the hopping, the interactions and the chemical potential $\mu$ or equivalently the average filling $\bar{n}$ depending if the grand canonical or canonical ensemble is considered. In what follows, we rely on the average filling where for $\bar{n}\in\mathbb{N}^{*}$, the SF-MI phase transition is of the Berezinskii-Kosterlitz-Thouless type at the critical value $(U/J)_{\mathrm{c}}\simeq 3.3$ for $\bar{n}=1$ [64, 65, 66, 67]. For non-integer fillings, the quantum system remains in the SF phase for any value of the dimensionless interaction parameter $U/J$.
@@
The dissipative quench dynamics associated to the BH chain is fully characterised by the Lindblad master equation which reads as:
$\frac{\mathrm{d}}{\mathrm{d}t}\hat{\rho}(t)=-i\left[\hat{H},\hat{\rho}(t)\right]+\sum_{R}\hat{L}_{R}\hat{\rho}(t)\hat{L}^{{\dagger}}_{R}-\frac{1}{2}\left\{\hat{L}^{{\dagger}}_{R}\hat{L}_{R},\hat{\rho}(t)\right\},$ (2)
@@
where $\hat{H}$ refers to the Hamiltonian of the BH chain defined at Eq. (1). $\hat{\rho}(t)$ refers to the time-dependent density matrix and $\hat{L}_{R}$ to the Lindblad jump operator acting on the lattice site $R$. For on-site two-body losses, $\hat{L}_{R}=\sqrt{\gamma}\hat{b}_{R}^{2}$ where $\gamma$ corresponds to the dissipation strength also called dissipation rate.
@@
To drive the BH chain out of equilibrium, the following quench procedure is considered. We start from an initial many-body quantum state corresponding to the ground state of the 1D BH model without dissipation implying $\gamma=0$. Initially, the density or equivalently the filling since $a=1$ is unitary, i.e. $\bar{n}=1$. Then, we let the quantum lattice model evolve with a non-zero value of the dissipation strength $\gamma>0$. The latter defines the first quench corresponding to an dissipation quench. The second one is performed on the dimensionless interaction parameter $U/J$ hence corresponding to an interaction quench. It is done by considering a pre-quench value $U_{\mathrm{i}}/J$ quenched to a post-quench value $U_{\mathrm{f}}/J$. Note that the hopping amplitude $J$ is fixed during the full real time evolution. The double quench dynamics is studied starting from a many-body ground state of the BH chain confined in the SF phase, i.e. $U/J<(U/J)_{\mathrm{c}}$. To characterise the system’s dynamics, we investigate a variety of quantum observables including $G_{1}(R,t)=\langle\hat{b}_{R}^{{\dagger}}(t)\hat{b}_{0}(t)\rangle_{c}$ the connected one-body correlation function and $G_{2}(R,t)=\langle\hat{n}_{R}(t)\hat{n}_{0}(t)\rangle_{c}$ the connected density-density correlation function. The phase and density fluctuations determined by $G_{1}$ and $G_{2}$ can be measured in ultracold-atom experiments using time-of-flight and fluorescence microscopy imaging respectively  [3, 36, 46, 47].
@@
To investigate theoretically the dissipative quench dynamics induced by on-site two-body losses of the BH chain initially confined in the SF-mean-field regime, we relied on an analytical result from the scientific publication of J.Despres et al. at Ref. [59]. In the latter paper, the set of equations of motion (EoMs) associated to the correlators $G_{\mathbf{k}}(t)=\langle\hat{b}^{{\dagger}}_{\mathbf{k}}\hat{b}_{\mathbf{k}}\rangle_{t}=\langle\hat{n}_{\mathbf{k}}\rangle_{t}$ and $F_{\mathbf{k}}(t)=\langle\hat{b}_{-\mathbf{k}}\hat{b}_{\mathbf{k}}\rangle_{t}$ for the BH model on a $D$-dimensional hypercubic lattice for long-range two-body losses has been calculated. Using the Lindblad master equation, the EoMs are given by:
@@
$\frac{\mathrm{d}}{\mathrm{d}t}F_{\mathbf{k}}(t)=i\left\langle\left[\hat{H}(t),\hat{b}_{-\mathbf{k}}\hat{b}_{\mathbf{k}}\right]\right\rangle_{t}+\sum_{\mathbf{R},\mathbf{R}^{\prime}}\langle\hat{L}_{\mathbf{R},\mathbf{R}^{\prime}}^{{\dagger}}\hat{b}_{-\mathbf{k}}\hat{b}_{\mathbf{k}}\hat{L}_{\mathbf{R},\mathbf{R}^{\prime}}\rangle_{t}-\frac{1}{2}\langle\hat{b}_{-\mathbf{k}}\hat{b}_{\mathbf{k}}\hat{L}_{\mathbf{R},\mathbf{R}^{\prime}}^{{\dagger}}\hat{L}_{\mathbf{R},\mathbf{R}^{\prime}}\rangle_{t}-\frac{1}{2}\langle\hat{L}_{\mathbf{R},\mathbf{R}^{\prime}}^{{\dagger}}\hat{L}_{\mathbf{R},\mathbf{R}^{\prime}}\hat{b}_{-\mathbf{k}}\hat{b}_{\mathbf{k}}\rangle_{t},$ (3b)
where $\hat{H}(t)$ denotes the quadratic form in Fourier space of the BH model in the SF-mean-field regime with a time-dependent filling $\bar{n}(t)$ or equivalently a time-dependent density $n(t)$. $\hat{b}_{\mathbf{k}}$ and $\hat{b}_{\mathbf{k}}^{{\dagger}}$ correspond to the bosonic annihilation and creation operators acting in the reciprocal space on the momentum $\mathbf{k}$; $\hat{n}_{\mathbf{k}}=\hat{b}^{{\dagger}}_{\mathbf{k}}\hat{b}_{\mathbf{k}}$ denotes the occupation number operator in Fourier space for the momentum $\mathbf{k}$. The previous set of equations is characterized by the Lindblad jump operator $\hat{L}_{\mathbf{R},\mathbf{R}^{\prime}}$ as well as the long-range dissipation strength $\gamma_{||\mathbf{R}-\mathbf{R^{\prime}}||}$ which read as:
$\hat{L}_{\mathbf{R},\mathbf{R}^{\prime}}=\sqrt{\gamma_{||\mathbf{R}-\mathbf{R}^{\prime}||}}\hat{b}_{\mathbf{R}}\hat{b}_{\mathbf{R}^{\prime}},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \gamma_{||\mathbf{R}-\mathbf{R}^{\prime}||}=\frac{\Gamma}{(1+||\mathbf{R}-\mathbf{R}^{\prime}||)^{\alpha}},$ (4)

### v2

# Quench Spectroscopy for Dissipative and (Non)-Hermitian Quantum Lattice Models

@@ Abstract
We extend the quench spectroscopy method to dissipative and isolated non-Hermitian quantum lattice models via the case study of the open Bose-Hubbard chain and the non-Hermitian transverse-field Ising chain respectively. We first investigate theoretically the dynamics of the open Bose-Hubbard chain confined in the superfluid phase induced by a sudden global quench on the dissipations and the repulsive interactions using the equation-of-motion approach. Using the same analytical approach, we then discuss the applicability of the quench spectroscopy to non-Hermitian quantum lattice models by considering the sudden global quench dynamics of the non-Hermitian transverse-field Ising chain confined in the paramagnetic phase. We finally generalize this spectroscopy method to isolated Hermitian quantum lattice models characterized by a quadratic fermionic or bosonic Hamiltonian. For this purpose, we consider the case study of the Hermitian version of the latter one-dimensional lattice model. The investigation is performed analytically for the bosonic and fermionic reformulations while considering for each case the equation-of-motion and quasiparticle theoretical approaches.
I Introduction
In the last decades, the major progress in the experimental control of quantum lattice models has given consequent momentum to the analytical and experimental investigation of the quench dynamics of isolated Hermitian quantum systems jurcevic2014 ; richerme2014 ; cheneau2012 ; polkovnikov2011 ; gogolin2016 ; calabrese2005 . Recently, a new spectroscopy method has been introduced theoretically and verified numerically using tensor-network-based techniques for isolated quantum spin and bosonic lattice models. The latter, justified using an eigenstate decomposition, is based on the dynamical properties of the model induced by weak sudden global or local quenches sanchezpalencia2019 ; villa2020 and is referred to as quench spectroscopy. This new method to determine low-lying excitation spectra has been recently applied experimentally to the dipolar XY model using neutral Rydberg atoms browaeys2024 and has provided promising results. However, since quantum systems cannot be purely isolated from their environment, it is natural that the study of open quantum systems has attracted a lot of interest in the last few years zoller2008 ; buchler2008 ; cirac2009 by taking into account dissipative effects such as local dephasing noise esposito2005 ; gaspard2005 ; Eisler_2011 ; kollath2018 ; turkeshi2021diffusion , incoherent hopping alba2020 and local gain/loss processes  langen2016 ; rauer2016 ; syassen2008 ; bo2013 ; ott2013 ; alba2022 ; rosso2021 ; rosso2023 ; mazza2023dissipative . These open quantum lattice models are governed by the well-known Lindblad master equation  fazio2024 ; breuer2007 ; ashida2020 ; daley2014 characterizing the time evolution of the quantum system density matrix. However, the possibility to apply the quench spectroscopy to open and non-Hermitian quantum lattice models has not been investigated and corresponds to the central point of this research work.
Experimentally, quantum simulators based on trapped ultra-cold atoms loaded in an artificial optical lattice generated by the interference of counter-propagating laser beams permit to simulate quantum lattice models bloch2012 ; gross2017 . Using bosonic atoms, the Bose-Hubbard model can thus be engineered bakr2010 ; chen2011 . Its out-of-equilibrium properties are accessible via quantum quenches realized experimentally by suddenly modifying the intensity of the laser beams controlling the artificial lattice depth and thus the ratio between the hopping amplitude and the two-body repulsive interaction strength cheneau2012 ; trotzky2012 . This tuning of the interaction ratio offers the possibility to scan the quantum phase transition between the gapless superfluid and gapped Mott-insulating phases for commensurate fillings of the lattice. Loss processes, inducing a rich and new physics such as the quantum Zeno effect schiro2022 , the loss-assisted quantum control dong2022 or the loss-induced cooling effect kohl2024 , can be generated using quantum simulators based on ultra-cold bosonic atoms where different number of particles can be involved. For instance, two-body losses can be engineered by light-assisted inelastic two-body collisions which can also occur naturally within this experimental platform aspect2000 ; franchi2017 ; tomita2017 ; syassen2008 ; weiner1999 ; bouganne2020 .
In this work, we propose to investigate the dynamics of the open Bose-Hubbard chain driven out of equilibrium via a sudden global quench on both the interaction and dissipation strengths. The dissipative effects consist here in on-site two-body losses. The term on-site refers to a loss process involving a single lattice site. This specific quantum lattice model as well as its quench dynamics can be simulated experimentally using ultra-cold bosonic atoms greiner2002b ; cheneau2012 ; langen2013 ; geiger2014 including a controllable strength of the on-site two-body losses engineered by light-assisted inelastic collisions tomita2017 ; syassen2008 ; weiner1999 ; bouganne2020 . The quench dynamics of the isolated Bose-Hubbard chain has already been extensively studied theoretically and numerically using for instance tensor network or variational Monte-Carlo techniques  calabrese2006 ; barmettler2012 ; kollath2007 ; moeckel2008 ; manmana2009 ; roux2010 ; navez2010 ; carleo2014 ; krutitsky2014 ; despres2018 ; despres2019 ; sanchezpalencia2019 ; villa2020 . However, the quench dynamics associated to the latter quantum model in the presence of dissipations, corresponding here to local two-body losses, has been under-researched analytically and numerically despresnew ; liu2022 ; liu2024 ; nagao2024 . Most importantly, the possibility to perform quench spectroscopy to dissipative quantum lattice models has not been studied. The latter statement is also valid for non-Hermitian quantum lattice models. This exotic class of quantum systems appear naturally when studying the dynamics of dissipative quantum lattice models in the so-called no-click limit daley2014 . Consequently, we also propose here to investigate the quench spectroscopy of the non-Hermitian ($s=1/2$ short-range interacting) transverse-field Ising chain confined in the paramagnetic phase. Finally, we extend the quench spectroscopy method to isolated Hermitian quantum lattice models. More precisely, we show that the quench spectroscopy method is suitable and reliable for any quadratic quantum lattice model being bosonic or fermionic. To do so, the case study of the transverse-field Ising chain confined in the $z$-polarized phase is considered. Its quench dynamics induced by a sudden modification of the external field as well as its quench spectral function are investigated analytically using the bosonic and fermionic reformulations of its Hamiltonian in the latter gapped quantum phase while comparing the equation-of-motion and quasiparticle theoretical approaches. The previous statements underline the importance of the research work presented here together with the possibility to confirm our conclusions experimentally; at least for dissipative and isolated Hermitian quantum lattice models.
@@ I Introduction
This article is organized as follows: in Sec. II, we start by introducing the dissipative one-dimensional Bose-Hubbard model and the quench procedure in Subsec. II.1. In Subsec. II.2, we benchmark the equation-of-motion and quasiparticle approaches to unravel the main dynamical features of the isolated Bose-Hubbard chain. In Subsec. II.3, we discuss the dissipative and interaction quench dynamics of the Bose-Hubbard chain where the loss processes correspond here to on-site (local) two-body losses. In Subsec. II.4, we investigate theoretically the quench spectroscopy of the latter dissipative quantum lattice model. In Sec. III, we discuss the experimental feasibility as well as the limitations associated to the quench spectroscopy method. Then, in Sec. IV, the quench spectroscopy approach is extended to non-Hermitian quantum lattice models by considering the non-Hermitian transverse-field Ising chain confined in the paramagnetic phase. In Sec. V, we move on to the Hermitian version of the latter spin model to validate the quench spectroscopy method to any quadratic quantum lattice model being bosonic or fermionic. This model has the advantage to be diagonalized using a bosonic or fermionic reformulation discussed in Subsec. V.1 and Subsec. V.2 respectively. Finally in Sec. VI, we present our conclusions and possible extensions of the presented research work.
II Dissipative quench dynamics of the Bose-Hubbard chain induced by local two-body losses
II.1 Model and quench procedure
@@
where $\hat{b}_{R}$ and $\hat{b}_{R}^{{\dagger}}$ correspond to the bosonic annihilation and creation operators acting on the lattice site $R\in\mathbb{N}$, $\hat{n}_{R}=\hat{b}^{{\dagger}}_{R}\hat{b}_{R}$ denotes the local occupation number operator associated to the lattice site index $R$, $J>0$ corresponds to the hopping amplitude, $U>0$ is the on-site repulsive two-body interaction strength. At equilibrium and zero-temperature, the quantum phase diagram of the BH chain has been extensively studied sachdev2001 ; cazalilla2011 . The latter displays a gapless superfluid (SF) phase and a gapped Mott-insulating (MI) phase with the so-called Mott lobes, determined by the competition between the hopping, the interactions and the chemical potential $\mu$ or equivalently the average filling $\bar{n}$ depending if the grand canonical or canonical ensemble is considered. In what follows, we rely on the average filling where for $\bar{n}\in\mathbb{N}^{*}$, the SF-MI phase transition is of the Berezinskii-Kosterlitz-Thouless type at the critical value $(U/J)_{\mathrm{c}}\simeq 3.3$ for $\bar{n}=1$ kuhner2000one ; kashurnikov1996exact ; ejima2011 ; rombouts2006 . For non-integer fillings, the quantum system remains in the SF phase for any value of the dimensionless interaction parameter $U/J$.
@@
The dissipative quench dynamics associated to the BH chain is fully characterized by the Lindblad master equation which reads as:
$\frac{\mathrm{d}}{\mathrm{d}t}\hat{\rho}=-i\left[\hat{H},\hat{\rho}\right]+\sum_{R}\hat{L}_{R}\hat{\rho}\hat{L}^{{\dagger}}_{R}-\frac{1}{2}\left\{\hat{L}^{{\dagger}}_{R}\hat{L}_{R},\hat{\rho}\right\},$ (2)
@@
where $\hat{H}$ refers to the Hamiltonian of the BH chain defined at Eq. (1). $\hat{\rho}$ refers to the time-dependent density matrix and $\hat{L}_{R}$ to the Lindblad jump operator acting on the lattice site $R$. For on-site two-body losses, $\hat{L}_{R}=\sqrt{\gamma}\hat{b}_{R}^{2}$ where $\gamma$ corresponds to the dissipation strength also called dissipation rate.
@@
To drive the BH chain out of equilibrium, the following quench procedure is considered. We start from an initial many-body quantum state corresponding to the ground state of the 1D BH model without dissipation implying $\gamma=0$. Initially, the density or equivalently the filling since $a=1$ is unitary, i.e. $\bar{n}=1$. Then, we let the quantum lattice model evolve with a non-zero value of the dissipation strength $\gamma>0$. The latter defines the first quench corresponding to a dissipation quench. The second one is performed on the dimensionless interaction parameter $U/J$ hence corresponding to an interaction quench. It is done by considering a pre-quench value $U_{\mathrm{i}}/J$ quenched to a post-quench value $U_{\mathrm{f}}/J$. Note that the hopping amplitude $J$ is fixed during the full real time evolution. The double quench dynamics is studied starting from a many-body ground state of the BH chain confined in the SF phase, i.e. $U/J<(U/J)_{\mathrm{c}}$. To characterize the system’s dynamics, we investigate a variety of quantum observables including $G_{1}(R,t)=\langle\hat{b}_{R}^{{\dagger}}(t)\hat{b}_{0}(t)\rangle_{c}$ the connected one-body correlation function and $G_{2}(R,t)=\langle\hat{n}_{R}(t)\hat{n}_{0}(t)\rangle_{c}$ the connected density-density correlation function. The phase and density fluctuations determined by $G_{1}$ and $G_{2}$ can be measured in ultra-cold-atom experiments using time-of-flight and fluorescence microscopy imaging respectively  cheneau2012 ; trotzky2012 ; langen2013 ; geiger2014 .
@@
To investigate theoretically the dissipative quench dynamics induced by on-site two-body losses of the BH chain initially confined in the SF-mean-field regime, we rely on an analytical result from Ref. despresnew . In the latter paper, the set of equations of motion (EoMs) associated to the correlators $G_{\mathbf{k}}(t)=\langle\hat{b}^{{\dagger}}_{\mathbf{k}}\hat{b}_{\mathbf{k}}\rangle_{t}=\langle\hat{n}_{\mathbf{k}}\rangle_{t}$ and $F_{\mathbf{k}}(t)=\langle\hat{b}_{-\mathbf{k}}\hat{b}_{\mathbf{k}}\rangle_{t}$ for the BH model on a $D$-dimensional hypercubic lattice for long-range two-body losses has been calculated. Using the Lindblad master equation, the EoMs are given by:
@@
$\frac{\mathrm{d}}{\mathrm{d}t}F_{\mathbf{k}}(t)=i\left\langle\left[\hat{H}(t),\hat{b}_{-\mathbf{k}}\hat{b}_{\mathbf{k}}\right]\right\rangle_{t}+\sum_{\mathbf{R},\mathbf{R}^{\prime}}\langle\hat{L}_{\mathbf{R},\mathbf{R}^{\prime}}^{{\dagger}}\hat{b}_{-\mathbf{k}}\hat{b}_{\mathbf{k}}\hat{L}_{\mathbf{R},\mathbf{R}^{\prime}}\rangle_{t}-\frac{1}{2}\langle\hat{b}_{-\mathbf{k}}\hat{b}_{\mathbf{k}}\hat{L}_{\mathbf{R},\mathbf{R}^{\prime}}^{{\dagger}}\hat{L}_{\mathbf{R},\mathbf{R}^{\prime}}\rangle_{t}-\frac{1}{2}\langle\hat{L}_{\mathbf{R},\mathbf{R}^{\prime}}^{{\dagger}}\hat{L}_{\mathbf{R},\mathbf{R}^{\prime}}\hat{b}_{-\mathbf{k}}\hat{b}_{\mathbf{k}}\rangle_{t};$ (3b)
$\hat{H}(t)=\frac{1}{2}\sum_{\mathbf{k}\neq\mathbf{0}}\mathcal{A}_{\mathbf{k}}(t)\left(\hat{b}^{{\dagger}}_{\mathbf{k}}\hat{b}_{\mathbf{k}}+\hat{b}_{-\mathbf{k}}\hat{b}^{{\dagger}}_{-\mathbf{k}}\right)+\mathcal{B}_{\mathbf{k}}(t)\left(\hat{b}^{{\dagger}}_{\mathbf{k}}\hat{b}^{{\dagger}}_{-\mathbf{k}}+\hat{b}_{\mathbf{k}}\hat{b}_{-\mathbf{k}}\right),$ (3c)
where $\hat{H}(t)$ denotes the quadratic form in the reciprocal Fourier space of the BH model in the SF-mean-field regime with a time-dependent filling $\bar{n}(t)$ or equivalently a time-dependent density $n(t)$. The latter is found using a standard mean-field approximation roux2013 ; despres2019 . The momentum- and time-dependent functions $\mathcal{A}_{\mathbf{k}}(t)$ and $\mathcal{B}_{\mathbf{k}}(t)$ depending on $n(t)$, or equivalently on $\bar{n}(t)$, are defined later on. $\hat{b}_{\mathbf{k}}$ and $\hat{b}_{\mathbf{k}}^{{\dagger}}$ correspond to the bosonic annihilation and creation operators acting in the reciprocal space on the momentum $\mathbf{k}$; $\hat{n}_{\mathbf{k}}=\hat{b}^{{\dagger}}_{\mathbf{k}}\hat{b}_{\mathbf{k}}$ denotes the occupation number operator in Fourier space for the momentum $\mathbf{k}$. The previous set of equations is characterized by the Lindblad jump operator $\hat{L}_{\mathbf{R},\mathbf{R}^{\prime}}$ as well as the long-range dissipation strength $\gamma_{||\mathbf{R}-\mathbf{R^{\prime}}||}$ which read as:
$\hat{L}_{\mathbf{R},\mathbf{R}^{\prime}}=\sqrt{\gamma_{||\mathbf{R}-\mathbf{R}^{\prime}||}}\hat{b}_{\mathbf{R}}\hat{b}_{\mathbf{R}^{\prime}},\leavevmode\nobreak\ \gamma_{||\mathbf{R}-\mathbf{R}^{\prime}||}=\frac{\Gamma}{(1+||\mathbf{R}-\mathbf{R}^{\prime}||)^{\alpha}},$ (4)